Part of a 6,000 investment made a 12% profit. The other part made an 8% loss. The net gain was 400. How much of the investment gained profit? ( answer is 4400 - show your solution )
Answer by lwsshak3(11628)

Part of a 6,000 investment made a 12% profit. The other part made an 8% loss. The net gain was 400. How much of the investment gained profit?
***
let x=Part of the investment that gained profit
let (6000-x)=Part of the investment showed loss
..
12%x-8%(6000-x)=400
.12x-480+.08x=400
.20x=880
x=4400
How much of the investment gained profit? $4400